OpenStudy (anonymous):

Using exact values, show that sin^2x-cos^2x= 1

OpenStudy (anonymous):

x=30 degrees

hartnn (hartnn):

but sin^2x-cos^2x is not equal to 1 sin^2x+cos^2x= 1

OpenStudy (anonymous):

nono you gotta use that for each angle. and the angle is x=30 degrees

OpenStudy (anonymous):

answer is (1/2)^2 + ( square root 3 / 3 ) ^2 = 1. idk how they got that

OpenStudy (anonymous):

*sqaure root 3 / 2

hartnn (hartnn):

u see the + sign in between?? thats because of sin^2x+cos^2x and not sin^2x-cos^2x

hartnn (hartnn):

rest values are correct,sin 30 =1/2 and cos 30 = (sqrt 3)/2

hartnn (hartnn):

which gives 1/4 + 3/4 =1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

damn hartnn went hard on ya lol

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):